OpenStudy (anonymous):

The length of a rectangle is 12 inches and needs to be changed to 8 inches. What is the scale factor of this dilation? F. 1/8 G. 1/3 H. 1/2 I. 2/3 (fan and medal) please explain how you got your answer(: ♥

OpenStudy (anonymous):

The answer is 2/3 12 x 2/3 = 8 You can also see it like this... divide 12 into 3 parts 1234......5678.......9,10,11,12 part one goes up to 4 part two goes up to 8 You only want 2 of the 3 parts which adds up to 8

OpenStudy (anonymous):

Oh! Thank you so much I'll be giving you your medal and fan(:

OpenStudy (anonymous):

Also if you divide 12/8 = 3/2